OpenStudy (anonymous):

Divide: -9 divided by (- 4) (show as an improper fraction)

OpenStudy (rane):

-9/4 = 2

hero (hero):

\[\frac{-9}{-4} = \frac{9}{4} = \frac{8 + 1}{4} = \frac{8}{4} + \frac{1}{4} = 2 \frac{1}{4}\]